h1p2:

Kay walks upstairs to make her legs stronger. Every day at work, she walks up 9 flights of stairs. If there are 13 steps in 1 flight of stairs, how many steps does Kay climb in 5 days at work?

KAPALOTJAY:

its multiplication ill tell u that

jhonyy9:

use the rule of 3 simple and write in this way 13 steps in 1 flight of stairs x steps in 9 flights of stairs so cross multiplie and calcule the x value x = 9*13/1 = 9*13 = ? steps in 9 flights per one day so to get how many steps Kay walks in 5 days just multiplie this number of steps per one day by 5 and will get sure easy the right answer hope helped understandably